A manufacturer of a brand of designer jeans realizes that many retailers charge less than the suggested retail price of $40. A random sample of 20 retailers reveals that the mean and the standard deviation of the prices of the jeans are $32 and $2.50 respectively. Estimate with 90% confidence the mean retail price of the jeans.


1
Expert's answer
2021-03-16T07:35:38-0400

"c = 90 \;\% = 0.90 \\

\bar{x}=32 \\

σ = 2.50 \\

n = 20 \\

df = n-1 = 19 \\

α = \frac{1-c}{2} = 0.05"

"t_{\u03b1\/2} = 1.328" (from t Table)

The margin error is:

"E = t_{α/2} \times \frac{s}{\sqrt{n}} = 1.328 \times \frac{2.50}{\sqrt{20}} = 0.742"

The mean retail price of the jeans with 90% confidence = 32±0.74